# Categorical Unsupervised Variational Acoustic Clustering

### v1

@@ I Introduction
The use of classification and clustering algorithms that allow for specialized processing are crucial for hardware-constrained applications like hearing aids [1, 2]. However, clustering algorithms require labels which can be scarce, as is the case for urban acoustic scenes datasets [3, 4], where they usually represent abstractions, e.g., the place where each audio file was recorded [5]. To optimize specialized processing, we need an unsupervised clustering method based on relevant characteristics of the acoustic signal [6] that is able to process complex and overlapped audio data.
While traditional methods struggle to cluster high-dimensional audio signals [7], variational autoencoders are a feasible option as they can learn without the need for labels [8] and generally do not require a massive number of parameters, as observed for other generative approaches [9]. Within the scope of interest, variational autoencoders were previously applied for unsupervised image clustering [10, 11, 12], and recently for the unsupervised clustering of spoken digits [6].
We build upon [6] by modifying the unsupervised variational acoustic clustering (UVAC) model to a categorical UVAC, inspired by the generative semi-supervised model [13]. The proposed approach employs a categorical distribution, providing efficient clustering performance even for datasets with strongly overlapped data points, as urban acoustic scenes. This is done using a Gumbel-Softmax (GS) distribution [14] for categorical reparametrization, allowing for training via backpropagation, with the softmax temperature serving as a mechanism to tune clustering performance. The results show impressive unsupervised clustering performance for a spoken digits dataset – considered for comparison with the baseline [6] – and for two real-world datasets of urban acoustic scenes.
@@ II Variational inference
We consider a dataset $\mathbf{X}=\{\mathbf{x}^{(i)}\}^{N}_{i=1}$ with $N$ independent and identically distributed (i.i.d.) samples. Each observation $\mathbf{x}^{(i)}$ is considered to be generated by a class $k$ of a categorical latent variable $y$ and a continuous latent variable $\mathbf{z}$, similarly to the M2 model in [13]. The joint distribution is given by
@@
[…]
to obtain the true posterior, with parameters $\theta$ that can be obtained by maximizing the log-likelihood $\log p_{\theta}(\mathbf{x}^{(i)})$. However, (2) has no closed-form solution for most real-world problems [8].
@@
[…]
@@
In (7), $D_{KL}(\cdot||\cdot)$ is the Kullback-Leibler (KL) divergence between two distributions and $\mathbb{E}_{q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})}[\log p_{\theta}(\mathbf{x^{(i)}}|y,\mathbf{z})]$ is the reconstruction error. Additionally, we include a non-trainable Lagrangian variable $\lambda$ for smoother training, which should be tuned based on data. The complete model with generation (1) and inference (3) is shown in Figure 1. In this model, each encoder and decoder are represented as a neural network (NN). To sample from a categorical distribution, the model uses a Gumbel-Softmax distribution [14], which we detail in the following.
@@ II-A Sampling from a categorical distribution
The basis of the approach comes from the Gumbel-Max trick [15]. Consider a categorical distribution with class probabilities $[\pi_{1},\ ...,\pi_{K}]$, defined as
@@
where $g_{k}\sim-\log(-\log(u))$ are i.i.d. samples drawn from the Gumbel distribution with $u\sim\mathcal{U}(0,1)$. Thereby, sampling from a discrete distribution is reduced by applying noise to a deterministic function. However, the $\operatorname*{\arg\max}$ operator is not differentiable. To address this issue, we change the $\arg\max$ operator by its differentiable approximation, the softmax function, and we obtain the Gumbel-Softmax distribution [14]:
@@ II-B Architecture
The categorical UVAC is shown in Figure 1. For the discriminative inference encoder $q_{\upsilon}(y|\mathbf{x}^{(i)})$, we consider a NN denoted by $h$ with parameters $\upsilon$ that outputs $K$ (soft) probabilities, $h_{\upsilon}(\mathbf{x}^{(i)})=[\pi_{1},\ ...,\pi_{K}]$. Moreover, the inference encoder $q_{\phi}(\mathbf{z}|y,\mathbf{x}^{(i)})$ is represented by a NN denoted by $f$ with parameters $\phi$ such that $f_{\phi}(y,\mathbf{x})=[\boldsymbol{\mu}_{\phi},\boldsymbol{\Sigma}_{\phi}]$, which outputs are used for sampling $\mathbf{z}$ via reparametrization trick [8] as $\mathbf{z}=\boldsymbol{\mu}_{\phi}+\boldsymbol{\Sigma}_{\phi}^{1/2}\boldsymbol{\epsilon}$, with an auxiliary random variable $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The generative decoder is given by a NN denoted by $g$ with parameters $\theta$ that reconstructs data $\mathbf{x}^{(i)}$, as $g_{\theta}(y,\mathbf{z})=[\hat{\mathbf{x}}^{(i)}]$. The NNs follow a similar architecture to that of [6], a state-of-the-art convolutional-recurrent autoencoder approach [16, 17], described as follows.
@@ III-A1 Spoken digit recognition
spoken digit recognition consists of identifying which digit was spoken in an audio utterance. The main acoustic feature is the digit itself, while other sound characteristics are minor features. For such a task, we use the AudioMNIST dataset [18] with 30000 audio samples, being 24000 randomly chosen for training, 3000 for validation, and 3000 for testing. Each file contains the recording of a spoken digit in english. We use the same pre-processing of data as described in [6], feeding a 1-second long magnitude spectrogram with trimmed frequency bins to the NNs.
@@ III-A2 Urban acoustic scene classification
we devise the classification of urban acoustic scenes using real-world datasets. Such a task is much more challenging as the (background) sound features resemble noise and greatly overlap in time and frequency. We consider two different datasets: TAU2019 [5], with 1200 audio recording files of 10 seconds from different acoustic scenes; and the UrbanSound8K (US8K) [19] dataset, with 8732 labeled sound excerpts of 10 different acoustic scenes, mostly 4-seconds long. Differently from AudioMNIST, we now use a mel-frequency cepstrum to reduce the input dimension without limiting frequency range, as acoustic scene classification can benefit from the broader frequency range.
We resample data to 16 kHz. The US8K audio files are zero-pad to four seconds, where we also employ a voice activity detection mask for the calculation of the reconstruction error, avoiding the network to cluster zero-padding [6]. We obtain a short-term Fourier transform (STFT) of 960 samples with a Hann window the same size and 50% overlap. Lastly, we apply a mel-frequency scale with 128 bins. The cepstrum is normalized by its mean and variance, and a min-max normalization to limit values from 0 to 1. We feed the NNs with a time-context window [20] of 4 seconds for the US8K dataset and 10 seconds for the TAU2019 dataset, which are the maximum duration of the dataset’s files.
Differently from AudioMNIST case, urban scene classification has direct application for hearing aids [1, 2], where different acoustic scenes result in different processing, which is proportional to the constraints of the device. We take two cases into account: the same number of clusters as labels in the dataset; and a reduced number of clusters. Specifically for the considered urban acoustic scene datasets, we consider 10 clusters for the higher end, as it matches the number of labels in the dataset. For the lower end, we take 5 clusters into account, as it is a significant reduction from 10, merging similar clusters, but still sufficient for effectively calculating clustering metrics. In practice, a higher number of clusters results in a more complex and “specialized” processing, representing a higher-end version of a hearing aid device. On the other hand, the lower cluster number could represent a more affordable version of the same device.
@@
TABLE I: Clustering metrics on the test set of the mentioned datasets, either considering labels as clusters or by applying K-means, UVAC and categorical (Cat.) UVAC as clustering methods, averaged over 10 independent runs. *results from [6].
Dataset Clusters Method Accuracy (%) $\uparrow$ NMI $\uparrow$ Silhouette $\uparrow$ DBI $\downarrow$ CHI ×$\times 10^{3}$uparrow
@@ III-B Metrics
Five metrics are considered: i) Unsupervised accuracy – the Hungarian algorithm [21] is used to match cluster labels to truth labels. The matched labels are used for calculating accuracy, ranging from 0 to 100%; ii) Normalized mutual information (NMI) – evaluates how much information is shared between cluster and truth labels [22], ranging from 0 to 1, being proportional to accuracy; iii) Silhouette score – measures how similar a data point is to its cluster in comparison to different clusters [23]. The value reflects on cohesion – how similar data points are in a cluster – and separation – how separate the clusters are. It ranges from -1 (bad) to 1 (good); iv) Davies-Bouldin index (DBI) – the average similarity rate of each cluster with its most similar cluster [24], being an indication of compactness and separation of clusters. The DBI score ranges from 0 (good) to infinity (bad); v) Caliński-Harabasz index (CHI) – measures the ratio of the sum of between-cluster to within-cluster dispersion [25] – distinctiveness and compactness. It ranges from 0 (bad) to infinity (good). The CHI is added on top of the metrics considered in [6] for a broader analysis. Accuracy and NMI tell us how predicted clusters deviate from labels, and the others show how good the clusters are.
@@ III-C Baselines
As a main baseline, we consider the UVAC model from [6]. We apply it with the exact same architecture as defined in the original paper. We expect UVAC to have a good clustering performance for the better-defined dataset, AudioMNIST, as the main feature – the spoken digit – is free of noise and other overlapping audio content. However, since UVAC takes a Gaussian mixture model (GMM) for clustering, it should fail for noise-resembling data, as the GMM lacks a mechanism to enforce clustering behavior, like the temperature $\tau$ of the GS. Additionally, we consider a classical approach, the K-means algorithm [26], as a low-complexity baseline.
@@ III-D Hyperparameters
We train the model described in Section II-B for 500 epochs, separately for each considered dataset. All NN parameters $\theta$, $\upsilon$, and $\phi$ are jointly optimized by maximizing (7) with the Adam algorithm [27]. For all cases of the proposed categorical UVAC and the baseline UVAC model [6], the initial learning rate is of $5\times 10^{-4}$, with exponential decay until the last epoch, reaching $5\times 10^{-5}$. The Gumbel-Softmax temperature $\tau$ in (9) of the categorical UVAC is annealed from 1.0 to 0.5 over the epochs, providing a “more uniform” distribution at the beginning of training, avoiding local minima, and converging to an “almost categorical” representation at end of training, enforcing clustering behavior. The value of $\lambda$ in (7) is chosen experimentally and we try to keep it as low as possible to avoid disturbing the balance of the variational lower bound, while still avoiding the collapse – unintended merging – of clusters. For AudioMNIST, we use $\tau=0.5$, while we change it to $\tau=2.0$ for both TAU2019 and UrbanSound8K datasets.

### v2

@@ I Introduction
The use of classification and clustering algorithms that allow for specialized processing are crucial for hardware-constrained applications like hearing aids [yook2015environmentadaptive, lamarche2010adaptive]. However, clustering algorithms require labels which can be scarce, as is the case for urban acoustic scenes datasets [fiorio2023semisupervised, han2016semisupervised], where they usually represent abstractions, e.g., the place where each audio file was recorded [mesaros2018tau]. To optimize specialized processing, we need an unsupervised clustering method based on relevant characteristics of the acoustic signal [fiorio2025unsupervised] that is able to process complex and overlapped audio data.
While traditional methods struggle to cluster high-dimensional audio signals [foote1999overview], variational autoencoders are a feasible option as they can learn without the need for labels [kingma2014autoencoding] and generally do not require a massive number of parameters, as observed for other generative approaches [kaplan2020scalinglawsneurallanguage]. Within the scope of interest, variational autoencoders were previously applied for unsupervised image clustering [ugur2020variational, jiang2017variational, dilokthanakul2017deep], and recently for the unsupervised clustering of spoken digits [fiorio2025unsupervised].
We build upon [fiorio2025unsupervised] by modifying the unsupervised variational acoustic clustering (UVAC) model to a categorical UVAC, inspired by the generative semi-supervised model [kingma2014semisupervised]. The proposed approach employs a categorical distribution, providing efficient clustering performance even for datasets with strongly overlapped data points, as urban acoustic scenes. This is done using a Gumbel-Softmax (GS) distribution [jang2017categorical] for categorical reparametrization, allowing for training via backpropagation, with the softmax temperature serving as a mechanism to tune clustering performance. The results show impressive unsupervised clustering performance for a spoken digits dataset – considered for comparison with the baseline [fiorio2025unsupervised] – and for two real-world datasets of urban acoustic scenes.
@@ II Variational inference
We consider a dataset $\mathbf{X}=\{\mathbf{x}^{(i)}\}^{N}_{i=1}$ with $N$ independent and identically distributed (i.i.d.) samples. Each observation $\mathbf{x}^{(i)}$ is considered to be generated by a class $k$ of a categorical latent variable $y$ and a continuous latent variable $\mathbf{z}$, similarly to the M2 model in [kingma2014semisupervised]. The joint distribution is given by
@@
$p_{\theta}(\mathbf{x}^{(i)})=\sum_{y}\int p_{\theta}(\mathbf{x^{(i)}},y,\mathbf{z})d\mathbf{z}$ (2)
to obtain the true posterior, with parameters $\theta$ that can be obtained by maximizing the log-likelihood $\log p_{\theta}(\mathbf{x}^{(i)})$. However, (2) has no closed-form solution for most real-world problems [kingma2014autoencoding].
@@
$\log p_{\theta}(\mathbf{x}^{(i)})=\log\sum_{y}\int q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})\frac{p_{\theta}(\mathbf{x^{(i)}},y,\mathbf{z})}{q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})}d\mathbf{z},$ (4)
where $\sum_{y}\int q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})[\cdot]d\mathbf{z}$ represents the expectation over $q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})$ applied to $\cdot$], i.e., $\mathbb{E}_{q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})}[\cdot]$. Rewriting (4) with the expectation operator gives us
@@
In (7), $D_{KL}(\cdot||\cdot)$ is the Kullback-Leibler (KL) divergence between two distributions and $\mathbb{E}_{q_{\upsilon,\phi}(y,\mathbf{z}|\mathbf{x}^{(i)})}[\log p_{\theta}(\mathbf{x^{(i)}}|y,\mathbf{z})]$ is the reconstruction error. Additionally, we include a non-trainable Lagrangian variable $\lambda$ for smoother training, which should be tuned based on data. The complete model with generation (1) and inference (3) is shown in Figure 1. In this model, each encoder and decoder are represented as a neural network (NN). To sample from a categorical distribution, the model uses a Gumbel-Softmax distribution [jang2017categorical], which we detail in the following.
@@ II-A Sampling from a categorical distribution
The basis of the approach comes from the Gumbel-Max trick [maddison2015asampling]. Consider a categorical distribution with class probabilities $\pi_{1},\ ...,\pi_{K}]$, defined as
@@
where $g_{k}\sim-\log(-\log(u))$ are i.i.d. samples drawn from the Gumbel distribution with $u\sim\mathcal{U}(0,1)$. Thereby, sampling from a discrete distribution is reduced by applying noise to a deterministic function. However, the $\operatorname*{\arg\max}$ operator is not differentiable. To address this issue, we change the $\arg\max$ operator by its differentiable approximation, the softmax function, and we obtain the Gumbel-Softmax distribution [jang2017categorical]:
@@ II-B Architecture
The categorical UVAC is shown in Figure 1. For the discriminative inference encoder $q_{\upsilon}(y|\mathbf{x}^{(i)})$, we consider a NN denoted by $h$ with parameters $\upsilon$ that outputs $K$ (soft) probabilities, $h_{\upsilon}(\mathbf{x}^{(i)})=[\pi_{1},\ ...,\pi_{K}]$. Moreover, the inference encoder $q_{\phi}(\mathbf{z}|y,\mathbf{x}^{(i)})$ is represented by a NN denoted by $f$ with parameters $\phi$ such that $f_{\phi}(y,\mathbf{x})=[\boldsymbol{\mu}_{\phi},\boldsymbol{\Sigma}_{\phi}]$, which outputs are used for sampling $\mathbf{z}$ via reparametrization trick [kingma2014autoencoding] as $\mathbf{z}=\boldsymbol{\mu}_{\phi}+\boldsymbol{\Sigma}_{\phi}^{1/2}\boldsymbol{\epsilon}$, with an auxiliary random variable $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The generative decoder is given by a NN denoted by $g$ with parameters $\theta$ that reconstructs data $\mathbf{x}^{(i)}$, as $g_{\theta}(y,\mathbf{z})=[\hat{\mathbf{x}}^{(i)}]$. The NNs follow a similar architecture to that of [fiorio2025unsupervised], a state-of-the-art convolutional-recurrent autoencoder approach [shaug2024speechenhancement, zhao2018convolutionalrecurrent], described as follows.
@@ III-A1 Spoken digit recognition
spoken digit recognition consists of identifying which digit was spoken in an audio utterance. The main acoustic feature is the digit itself, while other sound characteristics are minor features. For such a task, we use the AudioMNIST dataset [becker2023audiomnist] with 30000 audio samples, being 24000 randomly chosen for training, 3000 for validation, and 3000 for testing. Each file contains the recording of a spoken digit in english. We use the same pre-processing of data as described in [fiorio2025unsupervised], feeding a 1-second long magnitude spectrogram with trimmed frequency bins to the NNs.
@@ III-A2 Urban acoustic scene classification
we devise the classification of urban acoustic scenes using real-world datasets. Such a task is much more challenging as the (background) sound features resemble noise and greatly overlap in time and frequency. We consider two different datasets: TAU2019 [mesaros2018tau], with 1200 audio recording files of 10 seconds from different acoustic scenes; and the UrbanSound8K (US8K) [salamon2014us8k] dataset, with 8732 labeled sound excerpts of 10 different acoustic scenes, mostly 4-seconds long. Differently from AudioMNIST, we now use a mel-frequency cepstrum to reduce the input dimension without limiting frequency range, as acoustic scene classification can benefit from the broader frequency range.
We resample data to 16 kHz. The US8K audio files are zero-pad to four seconds, where we also employ a voice activity detection mask for the calculation of the reconstruction error, avoiding the network to cluster zero-padding [fiorio2025unsupervised]. We obtain a short-term Fourier transform (STFT) of 960 samples with a Hann window the same size and 50% overlap. Lastly, we apply a mel-frequency scale with 128 bins. The cepstrum is normalized by its mean and variance, and a min-max normalization to limit values from 0 to 1. We feed the NNs with a time-context window [fiorio2024spectral] of 4 seconds for the US8K dataset and 10 seconds for the TAU2019 dataset, which are the maximum duration of the dataset’s files.
Differently from AudioMNIST case, urban scene classification has direct application for hearing aids [yook2015environmentadaptive, lamarche2010adaptive], where different acoustic scenes result in different processing, which is proportional to the constraints of the device. We take two cases into account: the same number of clusters as labels in the dataset; and a reduced number of clusters. Specifically for the considered urban acoustic scene datasets, we consider 10 clusters for the higher end, as it matches the number of labels in the dataset. For the lower end, we take 5 clusters into account, as it is a significant reduction from 10, merging similar clusters, but still sufficient for effectively calculating clustering metrics. In practice, a higher number of clusters results in a more complex and “specialized” processing, representing a higher-end version of a hearing aid device. On the other hand, the lower cluster number could represent a more affordable version of the same device.
@@
TABLE I: Clustering metrics on the test set of the mentioned datasets, either considering labels as clusters or by applying K-means, UVAC and categorical (Cat.) UVAC as clustering methods, averaged over 10 independent runs. *results from [fiorio2025unsupervised].
Dataset Clusters Method Accuracy (%) $\uparrow$ NMI $\uparrow$ Silhouette $\uparrow$ DBI $\downarrow$ CHI $\times 10^{3}$ $\uparrow$
@@ III-B Metrics
Five metrics are considered: i) Unsupervised accuracy – the Hungarian algorithm [kuhn1955hungarian] is used to match cluster labels to truth labels. The matched labels are used for calculating accuracy, ranging from 0 to 100%; ii) Normalized mutual information (NMI) – evaluates how much information is shared between cluster and truth labels [vinh2010information], ranging from 0 to 1, being proportional to accuracy; iii) Silhouette score – measures how similar a data point is to its cluster in comparison to different clusters [rousseeuw1987silhouettes]. The value reflects on cohesion – how similar data points are in a cluster – and separation – how separate the clusters are. It ranges from -1 (bad) to 1 (good); iv) Davies-Bouldin index (DBI) – the average similarity rate of each cluster with its most similar cluster [davies1979cluster], being an indication of compactness and separation of clusters. The DBI score ranges from 0 (good) to infinity (bad); v) Caliński-Harabasz index (CHI) – measures the ratio of the sum of between-cluster to within-cluster dispersion [calinski1974dendrite] – distinctiveness and compactness. It ranges from 0 (bad) to infinity (good). The CHI is added on top of the metrics considered in [fiorio2025unsupervised] for a broader analysis. Accuracy and NMI tell us how predicted clusters deviate from labels, and the others show how good the clusters are.
@@ III-C Baselines
As a main baseline, we consider the UVAC model from [fiorio2025unsupervised]. We apply it with the exact same architecture as defined in the original paper. We expect UVAC to have a good clustering performance for the better-defined dataset, AudioMNIST, as the main feature – the spoken digit – is free of noise and other overlapping audio content. However, since UVAC takes a Gaussian mixture model (GMM) for clustering, it should fail for noise-resembling data, as the GMM lacks a mechanism to enforce clustering behavior, like the temperature $\tau$ of the GS. Additionally, we consider a classical approach, the K-means algorithm [hartigan1979kmeans], as a low-complexity baseline.
@@ III-D Hyperparameters
We train the model described in Section II-B for 500 epochs, separately for each considered dataset. All NN parameters $\theta$, $\upsilon$, and $\phi$ are jointly optimized by maximizing (7) with the Adam algorithm [kingma2015adam]. For all cases of the proposed categorical UVAC and the baseline UVAC model [fiorio2025unsupervised], the initial learning rate is of $5\times 10^{-4}$, with exponential decay until the last epoch, reaching $5\times 10^{-5}$. The Gumbel-Softmax temperature $\tau$ in (9) of the categorical UVAC is annealed from 1.0 to 0.5 over the epochs, providing a “more uniform” distribution at the beginning of training, avoiding local minima, and converging to an “almost categorical” representation at end of training, enforcing clustering behavior. The value of $\lambda$ in (7) is chosen experimentally and we try to keep it as low as possible to avoid disturbing the balance of the variational lower bound, while still avoiding the collapse – unintended merging – of clusters. For AudioMNIST, we use $\tau=0.5$, while we change it to $\tau=2.0$ for both TAU2019 and UrbanSound8K datasets.